OpenStudy (18jonea):

Which sample size will produce a margin of error of ±6.3%? 152 114 252 550

OpenStudy (18jonea):

@phi

OpenStudy (18jonea):

@robtobey

OpenStudy (18jonea):

@Directrix

OpenStudy (18jonea):

@Loser66

OpenStudy (ulink):

@Mehek14 @MissSmartiez

OpenStudy (ulink):

They'll help @18jonea. I'm not good at mathematics.

OpenStudy (anonymous):

The following may help: http://www.dummies.com/how-to/content/how-sample-size-affects-the-margin-of-error.html

OpenStudy (18jonea):

Can you help me work thrugh it @robtobey

OpenStudy (18jonea):

@Astrophysics

OpenStudy (ulink):

@MissSmartiez

OpenStudy (18jonea):

@phi

OpenStudy (18jonea):

@Mehek14

OpenStudy (18jonea):

@Kainui

OpenStudy (phi):

we need more info to do this

OpenStudy (18jonea):

like what this is all i got

OpenStudy (18jonea):

@phi

OpenStudy (phi):

I half remember that you need a level of confidence.

OpenStudy (18jonea):

ok so how do i solve this @phi

OpenStudy (phi):

We need more info or some assumptions. otherwise, I'm totally in the dark.

OpenStudy (18jonea):

same

OpenStudy (18jonea):

i can show you my notes

OpenStudy (18jonea):

@phi

OpenStudy (phi):

that might help

OpenStudy (18jonea):

does that help

OpenStudy (phi):

Yes, that helps a lot. They are saying (I guess it's a "rule of thumb" ) that the margin of error is about 1 / sqr(n)

OpenStudy (phi):

to do the problem, first change 6.3% to a decimal . what is that ?

OpenStudy (18jonea):

.063

OpenStudy (phi):

now set that equal to 1/ sqr(n) \[ 0.063= \frac{1}{\sqrt{n}}\] and "solve for n" do you know how to do that ?

OpenStudy (18jonea):

no

OpenStudy (phi):

can you solve for x if you had \[ \frac{1}{x} = 2 \] ?

OpenStudy (18jonea):

.5

OpenStudy (18jonea):

i got 252 would that be right

OpenStudy (phi):

you could multiply both sides by x to get \[ \frac{x}{x} = 2 x \] x/x is 1, so that is \[ 1 = 2x\] then divide both sides by 2 \[ \frac{1}{2} = x\]

OpenStudy (phi):

for your problem, rename sqr(n) as x (temporarily) \[ \frac{1}{x} = 0.063\] can you solve for x ?

OpenStudy (18jonea):

what

OpenStudy (phi):

first multiply both sides by x

OpenStudy (phi):

it will make sense in a minute

OpenStudy (18jonea):

I got +- square root 15.87 n= 251.85 then round to 252

OpenStudy (phi):

yes, that looks good

OpenStudy (phi):

\[ \frac{1}{\sqrt{n}}= 0.063\] multiply both sides by sqr(n) \[ \frac{\sqrt{n}}{\sqrt{n}}= 0.063{\sqrt{n}} \\1=0.063{\sqrt{n}}\] now divide both sides by 0.063 \[ \frac{1}{0.063}= \sqrt{n} \] finally square both sides to get rid of the square root \[ \left( \frac{1}{0.063}\right)^2= n \]

OpenStudy (phi):

typing into google: ( 1/0.063)^2 = we get 251.952633 or 252 rounded to the nearest whole number

OpenStudy (18jonea):

When 876 voters were polled, 67% said they were voting yes on an initiative measure. Find the margin of error and an interval that is likely to contain the true population proportion. ±29.6%; between 37.4% and 96.6% ±3.4%; between 63.6% and 70.4% ±34%; between 33% and 100% ±3.0%; between 64.0% and 70.0% Can you help me with this one

OpenStudy (18jonea):

@phi

OpenStudy (phi):

do you have any more notes? I'm not sure if we use 876 or 67% of 876.

OpenStudy (phi):

although, looking at the answer choices, it looks like we use the entire number in the sample. in other words, the measure of error is 1/sqr(876) what do you get ?

OpenStudy (18jonea):

around .04

OpenStudy (phi):

you need more decimals

OpenStudy (phi):

then change to a percent by multiplying by 100

OpenStudy (18jonea):

0.03378686892

OpenStudy (18jonea):

around 3.4 % which is choice b

OpenStudy (18jonea):

@phi

OpenStudy (phi):

yes. and the 3.4% is the error in their 67% so the range is 67-3.4 to 67+3.4 or 63.6% to 70.4%

OpenStudy (18jonea):

so b is correct @phi

OpenStudy (phi):

yes